Let X be a discrete random variable. If Pr(X<10) = 1/8, and Pr(X<=10) = 3/16, then what is Pr(X=10)? Please specify your answer in decimal terms and round your answer to the nearest hundredth (e.g., enter 12 percent as 0.12).
Answer #1

Now, we know that

P (X = 10) = P(X\leq 10) - P(X<10). =3/16 - 1/8

=3/16 - 2/16

=1/16

=0.0625

=0.06